# Qwen-GUI-3B: A Lightweight Vision–Language Model for Cross-Resolution GUI Grounding

### v1

@@ Abstract
This paper introduces Qwen-GUI-3B, a lightweight Vision-Language Model (VLM) specifically designed for Graphical User Interface grounding tasks, achieving performance competitive with significantly larger models. Unlike large-scale VLMs ($>$7B parameters) that are computationally intensive and impractical for consumer-grade hardware, Qwen-GUI-3B delivers strong grounding accuracy while being fully trainable on a single GPU (RTX 4090). The model incorporates several key innovations: (i) combine cross-platform, multi-resolution dataset of 24K examples from diverse sources including mobile, desktop, and web GUI screenshots to effectively address data scarcity in high-resolution desktop environments; (ii) a two-stage fine-tuning strategy, where initial cross-platform training establishes robust GUI understanding, followed by specialized fine-tuning on high-resolution data to significantly enhance model adaptability; and (iii) data curation and redundancy reduction strategies, demonstrating that randomly sampling a smaller subset with reduced redundancy achieves performance comparable to larger datasets, emphasizing data diversity over sheer volume. Empirical evaluation on standard GUI grounding benchmarks—including ScreenSpot, ScreenSpot-v2, and the challenging ScreenSpot-Pro—highlights Qwen-GUI-3B’s exceptional accuracy, achieving 84.9% on ScreenSpot and 86.4% on ScreenSpot-v2, surpassing prior models under 4B parameters. Ablation studies validate the critical role of balanced sampling and two-stage fine-tuning in enhancing robustness, particularly in high-resolution desktop scenarios. The Qwen-GUI-3B model and related resources are available at: https://github.com/Han1018/Qwen-GUI-3B.
@@ 1 Introduction
Figure 1: Left: Overall ScreenSpot benchmark accuracy across model sizes. Right: Per-category accuracy on ScreenSpot, ScreenSpot-v2, and ScreenSpot-Pro.
Graphical User Interface (GUI) grounding—the task of locating the correct UI element on a screen given a natural language instruction—is a foundational capability for building intelligent GUI agents. While large-scale Vision-Language Models (VLMs) with over 7B parameters have achieved strong performance on this task [11, 4, 10], their computational requirements render them impractical for deployment on commodity hardware. This constraint has motivated recent interest in developing compact models that retain high grounding accuracy while remaining trainable and deployable on a single consumer GPU (e.g., RTX 4090) [8, 9].
@@
To address these issues, we propose Qwen-GUI-3B, a lightweight 3B-parameter vision-language model optimized for GUI grounding in cross-platform and resolution-diverse environments. Despite its compact size, the model achieves accuracy on par with 7B-scale baselines on standard benchmarks [4, 10, 11], demonstrating that model performance can be significantly improved through targeted training strategies and data selection rather than scale alone (see Figure 1).
@@ 3.3 Efficient Training Pipeline and Comparison to Prior Work
This work adopts a streamlined training pipeline that emphasizes simplicity, data efficiency, and computational accessibility. In contrast to recent GUI agent frameworks that introduce complex architectures or rely on large-scale data acquisition and reinforcement learning infrastructure, our approach demonstrates that a standard supervised fine-tuning setup—when paired with appropriate data and training strategy—is sufficient to achieve strong grounding performance.
@@
Compared to ShowUI [8], which also targets lightweight GUI grounding using a 2B model and 256K training examples, Qwen-GUI-3B further simplifies the pipeline by avoiding custom architectural changes. Whereas ShowUI introduced a UI-guided token pruning module to reduce visual redundancy, our model benefits implicitly from the resolution diversity of the dataset, allowing the base architecture to focus attention on relevant interface regions without external guidance.
The overall training process is kept minimal: fine-tuning a 3B backbone using LoRA-based adaptation [6], mixed-precision training, and gradient accumulation on a single RTX 4090 GPU. No auxiliary modules, external tools, or architectural modifications are introduced.
@@ 4.1 Experimental Setup and Evaluation
All experiments are conducted using a single NVIDIA RTX 4090 GPU (24,GB VRAM). Model training follows a resource-efficient setup based on DeepSpeed ZeRO-2 optimization and FlashAttention (SDPA) to reduce memory overhead and improve attention throughput. Fine-tuning is performed using LoRA with rank 8 and $\alpha=16$, applied to selected transformer layers [6]. Mixed-precision (FP16) training is used throughout, with a batch size of 1 and gradient accumulation steps set to 48. Each training epoch consists of 122 steps.
@@ 4.2 Main Results: Accuracy and Model Comparison
We evaluate Qwen-GUI-3B on three standard GUI grounding benchmarks: ScreenSpot [3], ScreenSpot-v2 [10], and the newly proposed ScreenSpot-Pro [7]. Despite its compact 3B parameter scale, the model achieves strong performance across all settings.
On ScreenSpot, Qwen-GUI-3B achieves an accuracy of 84.9%, and on the cleaned ScreenSpot-v2, it reaches 86.4%, setting a new benchmark among all sub-4B models. Compared to the previous best 2B-scale model, UI-TARS-2B [9], Qwen-GUI-3B yields a +2.6% gain on ScreenSpot and +1.7% on ScreenSpot-v2. These results indicate that our training strategy and dataset design effectively close the performance gap between small and mid-sized models. Beyond sub-4B models, Qwen-GUI-3B performs on par with several 7B-scale systems, such as Aguvis-7B [11] and OS-Atlas-7B [10]. These results suggest that, with proper adaptation and data curation, a 3B model can achieve competitive performance typically associated with models 2$\times$ its size. See Table 1 for detailed results.
On ScreenSpot-Pro, a challenging benchmark that includes professional software environments (e.g., Photoshop, AutoCAD) with dense visual layouts and smaller UI targets, Qwen-GUI-3B achieves a success rate of 28.7%. This surpasses the strongest prior 2B baseline (27.7%) and exceeds several 7B models, further validating its robustness in high-resolution, real-world applications. See Table 2 for details.
These results establish Qwen-GUI-3B as a strong candidate for GUI grounding under constrained resources. The model demonstrates that with a compact architecture and carefully designed training strategy, it is possible to attain competitive accuracy without relying on extensive model scale or massive datasets.
@@ 4.3 Ablation Study
Balanced Sampling. To address the inherent imbalance in GUI data across platforms, we compare uniform sampling (1:1:1 across mobile, web, and desktop) with joint training that reflects the original data distribution. As shown in Table 3, balanced sampling yields a modest improvement in overall accuracy (+0.9%) and a slight gain on desktop-specific performance (+1.3%). While the improvement is not large, it suggests that balancing platform representation during training can help stabilize performance on under-represented high-resolution interfaces [8, 2].
@@ 5 Conclusion
This work introduces Qwen-GUI-3B, a 3B-parameter vision-language model designed for GUI grounding under realistic resource constraints. Despite its compact size, the model delivers competitive accuracy across cross-platform benchmarks, including mobile, web, and desktop scenarios. These results demonstrate that lightweight models, when paired with carefully designed training strategies, can close the performance gap to significantly larger systems [9, 11, 10].
Our approach emphasizes data and training efficiency over scale. By integrating a compact yet diverse dataset (24K examples) and applying a two-stage fine-tuning strategy [4, 8], Qwen-GUI-3B achieves state-of-the-art performance among sub-4B models and rivals several 7B-scale baselines. Importantly, this is accomplished using a single consumer-grade GPU, without introducing architectural modifications or external modules [6].
@@
Qwen-GUI-3B provides a practical and reproducible strategy for building deployable GUI agents. This study affirms that model compactness need not compromise grounding performance, and that targeted design choices in data construction and training schedules play a critical role in maximizing capability within constrained computational budgets.

### v3

# ZonUI-3B: Competitive GUI Grounding with a 3B VLM Trained on a Single Consumer GPU

@@ Abstract
In this paper, we present ZonUI-3B, a lightweight Vision-Language Model (VLM) that can be fully trained on a single consumer-grade GPU (RTX 4090) while delivering performance comparable to significantly larger models on GUI grounding tasks. The model incorporates several key innovations: (i) combine cross-platform, multi-resolution dataset of 24K examples from diverse sources including mobile, desktop, and web GUI screenshots to effectively address data scarcity in high-resolution desktop environments; (ii) a two-stage fine-tuning strategy, where initial cross-platform training establishes robust GUI understanding, followed by specialized fine-tuning on high-resolution data to significantly enhance model adaptability; and (iii) data curation and redundancy reduction strategies, demonstrating that randomly sampling a smaller subset with reduced redundancy achieves performance comparable to larger datasets, emphasizing data diversity over sheer volume. Empirical evaluation on standard GUI grounding benchmarks—including ScreenSpot, ScreenSpot-v2, and the challenging ScreenSpot-Pro—highlights ZonUI-3B’s exceptional accuracy, achieving 84.9% on ScreenSpot and 86.4% on ScreenSpot-v2, surpassing prior models under 4B parameters. Ablation studies validate the critical role of balanced sampling and two-stage fine-tuning in enhancing robustness, particularly in high-resolution desktop scenarios. The ZonUI-3B model and related resources are available at: https://github.com/Han1018/ZonUI-3B.
@@ 1 Introduction
Figure 1: Overall ScreenSpot benchmark accuracy across model sizes.
Graphical User Interface (GUI) grounding—the task of locating the correct UI element on a screen given a natural language instruction—is a foundational capability for building intelligent GUI agents. While large-scale Vision-Language Models (VLMs) with over 7B parameters have achieved strong performance on this task [11, 4, 10], their training requirements are prohibitively expensive, making them inaccessible for many researchers and practitioners without high-end hardware. This has sparked growing interest in developing compact models that retain high grounding accuracy [8, 9].
@@
To address these issues, we propose ZonUI-3B, a lightweight 3B-parameter vision-language model optimized for GUI grounding in cross-platform and resolution-diverse environments. Despite its compact size, the model achieves accuracy on par with 7B-scale baselines on standard benchmarks [4, 10, 11], demonstrating that model performance can be significantly improved through targeted training strategies and data selection rather than scale alone (see Figures 1 and 2).
@@
Figure 2: Per-category accuracy on ScreenSpot, ScreenSpot-v2, and ScreenSpot-Pro.
@@ 3.3 Efficient Training Pipeline and Comparison to Prior Work
We present a streamlined training pipeline that emphasizes simplicity, data efficiency, and computational accessibility. In contrast to recent GUI agent frameworks that introduce complex architectures or rely on large-scale data acquisition and reinforcement learning infrastructure, our method demonstrates that a standard supervised fine-tuning setup—when paired with appropriate data and training strategy—is sufficient to achieve strong grounding performance.
@@
Compared to ShowUI [8], which also targets lightweight GUI grounding using a 2B model and 256K training examples, ZonUI-3B further simplifies the pipeline by avoiding custom architectural changes. Whereas ShowUI introduced a UI-guided token pruning module to reduce visual redundancy, our model benefits implicitly from the resolution diversity of the dataset, allowing the base architecture to focus attention on relevant interface regions without external guidance.
Our training process is kept minimal: fine-tuning a 3B backbone using LoRA-based adaptation [6] and gradient accumulation on a single RTX 4090 GPU. No auxiliary modules, external tools, or architectural modifications are introduced.
@@ 4.1 Experimental Setup and Evaluation
Our experiments are conducted using a single NVIDIA RTX 4090 GPU (24,GB VRAM). Model training follows a resource-efficient setup based on DeepSpeed ZeRO-2 optimization and FlashAttention (SDPA) to reduce memory overhead and improve attention throughput. Fine-tuning is performed using LoRA with rank 8 and $\alpha=16$, applied to selected transformer layers [6]. Mixed-precision (FP16) training is used throughout, with a batch size of 1 and gradient accumulation steps set to 48. Each training epoch consists of 122 steps.
@@ 4.2 Main Results: Accuracy and Model Comparison
We evaluate ZonUI-3B on three standard GUI grounding benchmarks: ScreenSpot [3], ScreenSpot-v2 [10], and the newly proposed ScreenSpot-Pro [7]. Despite its compact 3B parameter scale, the model achieves strong performance across all settings.
On ScreenSpot, ZonUI-3B achieves an accuracy of 84.9%, and on the cleaned ScreenSpot-v2, it reaches 86.4%, setting a new benchmark among all sub-4B models. Compared to the previous best 2B-scale model, UI-TARS-2B [9], ZonUI-3B yields a +2.6% gain on ScreenSpot and +1.7% on ScreenSpot-v2. These results indicate that our training strategy and dataset design effectively close the performance gap between small and mid-sized models. Beyond sub-4B models, ZonUI-3B performs on par with several 7B-scale systems, such as Aguvis-7B [11] and OS-Atlas-7B [10]. These results suggest that, with proper adaptation and data curation, a 3B model can achieve competitive performance typically associated with models 2$\times$ its size. See Table 1 for detailed results.
On ScreenSpot-Pro, a challenging benchmark that includes professional software environments (e.g., Photoshop, AutoCAD) with dense visual layouts and smaller UI targets, ZonUI-3B achieves a success rate of 28.7%. This surpasses the strongest prior 2B baseline (27.7%) and exceeds several 7B models, further validating its robustness in high-resolution, real-world applications. See Table 2 for details.
These results establish ZonUI-3B as a strong candidate for GUI grounding under constrained resources. The model demonstrates that with a compact architecture and carefully designed training strategy, it is possible to attain competitive accuracy without relying on extensive model scale or massive datasets.
@@ 4.3 Ablation Study
To address the inherent imbalance in GUI data across platforms, we compare uniform sampling (1:1:1 across mobile, web, and desktop) with joint training that reflects the original data distribution. As shown in Table 3, balanced sampling yields a modest improvement in overall accuracy (+0.9%) and a slight gain on desktop-specific performance (+1.3%). While the improvement is not large, it suggests that balancing platform representation during training can help stabilize performance on under-represented high-resolution interfaces [8, 2].
@@ 5 Conclusion
This work introduces ZonUI-3B, a 3B-parameter vision-language model designed for GUI grounding under realistic resource constraints. Despite its compact size, the model delivers competitive accuracy across cross-platform benchmarks, including mobile, web, and desktop scenarios. These results demonstrate that lightweight models, when paired with carefully designed training strategies, can close the performance gap to significantly larger systems [9, 11, 10].
Our approach emphasizes data and training efficiency over scale. By integrating a compact yet diverse dataset (24K examples) and applying a two-stage fine-tuning strategy [4, 8], ZonUI-3B achieves state-of-the-art performance among sub-4B models and rivals several 7B-scale baselines. Importantly, this is accomplished using a single consumer-grade GPU, without introducing architectural modifications or external modules [6].
@@
ZonUI-3B provides a practical and reproducible strategy for building deployable GUI agents. This study affirms that model compactness need not compromise grounding performance, and that targeted design choices in data construction and training schedules play a critical role in maximizing capability within constrained computational budgets.